OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x2 - 5x + 5 = 0

OpenStudy (johnweldon1993):

\[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (johnweldon1993):

You have your equation...and the 'a' 'b' and 'c' correspond to the coefficients of the equation a = -2 b = -5 c = 5 Plug those in and you will get your 2 results

OpenStudy (anonymous):

\[-\frac{ 5 }{ 4} \pm \frac{ \sqrt{65} }{ 4}\] I did that and I got this as an answer...I'm guessing it is correct?